OpenStudy (vera_ewing):

When 95.4 g of a metal absorbs 849 J of heat, the temperature of the metal increases from 25.0°C to 48.0°C. What is the specific heat of the metal? A. 0.387 J/g•°C B. 2.81 J/g•°C C. 5.39 J/g•°C D. - 2.81 J/g•°C

OpenStudy (anonymous):

Solve: \[\color{green}{c = \frac{E}{m \Delta T}}\] So, \[\color{green}{c = \frac{849~\text{J}}{(0.095\;4~\text{kg}) \times (48.0^{\circ}\text{C} - 25.0^{\circ}\text{C})}}\]

OpenStudy (vera_ewing):

@JoannaBlackwelder It's C right?

OpenStudy (joannablackwelder):

No, @ghuczek is correct. :)

OpenStudy (vera_ewing):

Ok thank you :) Can you help me with a few more?